if the first three terms of an increasing AP are the roots of the cubic 4x^3-24x^2+23x+18=0,then find S19.

Grade:11

1 Answers

Sunil Raikwar
9 years ago
let the roots of the equation a-d,a,a+d
sum of roots=6
3a=6
a=2
product of roots=-18/4
a(a2-d2)=-9/2
4-d2=-9/4
d2=25/4
d=5/2
S19=19/2{2x2+(19-1)5/2}
S19=19/2(49)=931/2